# Explainable Deep Learning Analysis for Raga Identification in Indian Art Music

### v1

@@ IV-A Feature Extraction
We use chromagram features, extracted from the audios using librosa library on python for our task. In our dataset, we have performances by various artists, each singing at different tonics. Tonic is just the reference note for construction of any Raga. Music performances in the same Raga can exhibit structural similarities when they use the same tonic. We performed tonic normalization on the extracted chromagram features, which aligns all songs to a common reference note, facilitating the classification model’s ability to discern Raga patterns irrespective of the original tonic used by the artists.
@@ IV-B Tonic Normalisation
For a given audio signal x, let the short time fourier transform is given by $X(n,k)$. The center frequency of a pitch in an equal-tempered midi scale is given by:
$F_{pitch}(p)=440\times 2^{(p-69)/12}$ (1)
From these, we get a pitch-based log-frequency spectrogram, which converts the whole spectrogram having frequency axis equal to NFFT to have 128 values corresponding to the 128 midi frequencies. We basically assign each of the $X(n,k)$ to the closest $f_{coef(k)}$ frequency. So, for each of the 128 pitch values $p\in[0:127]$, we have:
$P(p)={k:F_{pitch}(p-0.5)<=F_{coef}(k)<F_{pitch}(p+0.5)}$ (2)
So, we have the log frequency spectrogram $Y_{LF}$ given by:
$Y_{LF}(n,p)=\sum_{k\in P(p)}|X(n,k)|^{2}$ (3)
To obtain chromagram features, we simply sum up the pitch values which differ by some integer multiple of an octave as they are perceived to be the same note musically:
@@
Now, to shift the chromagram representation of any song by a factor of "$a"$ semitones, we are essentially looking to get a new $X^{\prime}(n,k)$ such that:
@@
Shifting by "$a"$ semitones essentially means multiplying the frequency values in a spectrogram by $2^{a/12}$, where "$a"$ is a discrete value such that $a\in[1:11]$. This range for "$a"$ is chosen because a value greater than 11 would result in an octave shift, which is not a consideration for us since it would return to the same musical note. Now, we will get log-frequency spectrogram as:
@@
Comparing Equation: 8 with 7, we can see that for a given $k^{\prime}$, which gets shifted ${}^{\prime}a^{\prime}$ semitones, the set of frequency values given by $P(p^{\prime})$ is just all the frequency values present ${}^{\prime}a^{\prime}$ chromagram bins higher. i.e., the pitch values, represented by $F_{\text{pitch}}(p)$, undergo a shift when transitioning from $k^{\prime}$ to $k$ or vice versa, where the difference between the two pitches is given by the factor ${}^{\prime}a^{\prime}$. This shift signifies a displacement of the frequency values associated with the musical notes, particularly within the context of a chromagram bin. It implies that the set of pitch values $P(p)$, forming a bin in the chromagram, undergoes a uniform displacement by a factor of ${}^{\prime}a^{\prime}$. Consequently, while the bins themselves remain unchanged, the pitch values contained within each bin experience a consistent shift, maintaining the relative positioning of musical notes while adjusting their absolute frequencies. So, in practice, for applying tonic normalisation to our audio data, we just shift the extracted chromagram bins through the shorter dimension in a cyclic manner.
@@ IV-C Classification Model Description
To determine the best performing model, we conduct an ablation study, experimenting with various CNN and LSTM-based architectures. Our first model, CN1+T, utilizes a simple CNN architecture with convolutional, max-pooling, and batch normalization layers, followed by a flatten layer, in a very similar way to a common Image classification model. Then the CN1+LSTM+T model extends this by replacing the flatten layer with an LSTM layer, ideal for processing time-series data like audio signals. Also, we provide tonic normalization as discussed in Section:IV-B and test the model performance. The CN2+LSTM model features a more complex CNN-LSTM architecture with 1D max-pooling to preserve temporal information, as shown in Fig: 2. Then the CN2+LSTM+T model just includes tonic normalization in the input features, following the exact configuration of CN2+LSTM model. The selection of layer configurations and activation functions is carefully optimized for classification performance. These models will be further detailed in sectionV, where we evaluate their effectiveness in ARI.
@@ IV-D Classification Model Training
We train our DL models for the 12-class classification task, while using softmax activation in the last layer. Let $z_{i}$ represents the output logit for the $i^{th}$ neuron in the last dense layer having 12 neurons. Softmax function applied to these logits converts them into probability values as:
$p_{i}=softmax(z_{i})=\frac{e^{z_{i}}}{\sum_{j=1}^{N}e^{z_{j}}}$ (9)
Here, $p_{i}$ is the probability score for $i^{th}$ neuron in the layer, $N=12$ as we have a 12-class classification problem. Also, $i\in[1:N]$, so we get 12 probability values, such that $\sum_{i}{p_{i}}=1$. Then for training, we use categorical cross-entropy loss function given by:
$L_{CCE}=\sum_{i=1}^{N}t_{i}log(p_{i})$ (10)
Here, $t_{i}$ is the $i^{th}$ true label. In this case, $t_{i}$ will be 1 for only 1 neuron and 0 for all other neurons corresponding to a particular input. The training is carried out for multiple epochs until the model performances converges measured by precision, recall and f-1 score. In our experiments, we use cross-validation to assess the robustness and generalization ability of our models for the classification task.
@@ IV-E Grad-CAM++
GradCAM++ builds upon GradCAM (Gradient-weighted Class Activation Maps), with a key advancement to it. Initially, we reload the CN2+LSTM+T model trained in above steps where the last CNN layer comprises 256 learned feature maps in our case as visible in the Fig: 2. To create a class-specific GradCAM++ activation map, each of these feature maps is weighted and then a weighted sum is taken to generate the final activation map or saliency map for the target class. We compute the weights of feature maps using the weights of those last CNN blocks, given by the equation:
@@
Here, $A^{k}_{ij}$ represents each feature map (so we have $k$ feature maps and $k\in[1:256]$). Also, we have $c$ classes with $c\in[1:12]$. $\alpha_{c}^{k}$ represents the weights of $k^{th}$feature map for $c^{th}$ class. $Z$ is the dimension of each feature map, and $y^{c}$ represents the output logits whose gradients we compute using the chain rule. Using these computed weights as the weights of each of the $k$ feature maps, we get the final Grad-CAM attribution map, given by:
@@
$L^{c}_{Grad-CAM}$ represents the final attribution map for GradCAM, which is obtained by summing all the feature maps up-sampled to the original input size, given by $A^{k}$, with each $A^{k}$ weighted with a corresponding weight $\alpha_{k}^{c}$ according to a given output prediction class $c$. After that, ReLU is applied to keep only the positively impacting salience regions and remove the regions with negative gradient in the output saliency map or attribution map. But, to obtain the activation map for GradCAM++, we weight each pixel separately rather than weighting the whole feature map with a single weight. The new weights are given by:
@@
Here, $A^{k}_{ij}$ represents each pixel in the $k^{th}$ feature map. This $w^{kc}_{ij}$ aims to weigh each pixel separately, rather than the whole activation map. So, this $w^{kc}_{ij}$ is given by:
@@
Using these, we finally obtain the attribution map for Grad-CAM++ given by:
@@ IV-F SoundLIME
SoundLIME[36] is specifically designed for audio, analogous to LIME’s[33] application in images. Here also, we need the pre-trained classifier model, but we treat it as a black-box model with no access to its gradients. Here, we temporally divide the chromagram into 30 equal parts or super-pixels, with one super-pixel representing the chromagram representation for each second of the audio. Next, we create 150 random perturbations for the given test input by applying random multi-hot masks of shape 30x1, meaning that each element of the mask either suppresses or retains each super-pixel. An exponential kernel is defined on L2 distance function $D(x,z)$ between the original image $x$ and the perturbed image $z$, denoted by $\pi_{x}(z)$:
@@ IV-G Saliency Evaluation Method
To evaluate the salient regions identified by the two ExAI models, we use the output of ExAI models, and compare them with the expert-annotated salient regions. We compute the precision for top 1 to 10 seconds for both the ExAI models. In our task, the concern lies solely with determining whether the top predicted regions are correct. Hence, precision is the best metric to employ in this situation, whereas recall and accuracy are not helpful. We also explain one correct and one incorrect prediction of the classifier and analyse their salient region. Also, the evaluation of correctness of ExAI model predictions is not required in our case, as we are solely focused on the salient regions’ correctness within the top few seconds in context of HCM, rather than assessing the ExAI model’s performance.
@@ V-A Data Pre-processing
To prepare the audio files for analysis, all non-musical speech has been removed. So, we firstly distribute the whole songs to train, test and validation sets, making sure that there is at least 1 song of each class in the test and validation sets. Then we split the songs into 30 second snippets for all the 3 splits with percentages of approximately 79%, 15%, and 6%, for training, validation, and test sets respectively as shown in Fig: 3 Notice that we split the dataset into train,test and validation sets first on file level, and then extract split them into 30-second clips. We do not directly create 30-seconds clips and then shuffle them as we have a possibility that we have snippets of the same song in train, test and validation set, which will add an additional bias in our dataset and will hinder the generalizability of our model.
@@
Also, the last part of each music recording is removed as it is mostly found to be having only 3 to 10 seconds of music, which may not even contain any salient region for getting classified to a particular raga and just serves the purpose of confusing the model. Also, generally in HCM performances, the endings of clips lasting 20 seconds or more may simply consist of a single note prolonged for a few seconds, which may not accurately represent the current Raga class. Thus, removing these endings should not adversely affect the analysis.
For our work, we extract chromagram features for these audio snippets, making each input sample of shape (938,12), having 938 time frames having 12 chromagram bins each.
All the songs, marked with their respective tonic by the music experts go through tonic normalisation as explained in IV-B. We make Note ”A” as the reference note for all and normalise accordingly. To achieve this, we just cyclically shift the obtained chromagram features through their shorter dimension throughout the 938 frames for each 30-sec file to obtain the desired normalisation.
@@ V-B Classification results
We chose audio files belonging to 12 classes having the most number of audio files out of the dataset and train a 12-class classifier using the models discussed in IV-C. Classification results for our Ablation studies are shown in table: II (all values rounded off to 2 decimal places)
TABLE II: Combined Classification Report for Different Models
@@ V-B1 CN1+T
This model achieved an overall weighted f1-score of 0.78 on the test data for all 12 classes after cross-validation. The maximum and minimum f1-scores during cross-validation were 0.80 and 0.73, respectively. Clearly, this architecture is not be suitable for audio data, as flattening does not appropriately handle the dependent framewise information required for Raga detection in case of such time-series data.
@@ V-B2 CN1+LSTM+T
This model showed improved classification performance with an f1-score of 0.83 after cross-validation. The maximum and minimum f1-score observed during cross-validation were 0.87 and 0.79, respectively. This model outperformed the CN1+T model, highlighting the importance of LSTM in stead of flattening in such scenarios.
@@ V-B3 CN2+LSTM
Despite multiple training attempts and hyper-parameter tunings, this model did not perform well on the training data itself, even with a similar architecture to our main model (CNT+LSTM+T) or a even deeper CNN-LSTM model. This underscores the importance of tonic normalization for the ARI task. We also experimented with removing classes with very low f1-scores, simplifying the problem. While this led to some improvement in training, it suggests a lack of generalization to any number of Raga classes using this approach.
@@ V-B4 CN2+LSTM+T
This model consistently achieved the highest performance among all models, with an f1-score of 0.89 after 6-fold cross-validation. For four classes, the CN1+LSTM+T model showed better results, however, the proposed model closely follows, demonstrating strong training across all 12 classes and better generalizability. The maximum and minimum f1-scores observed during cross-validation are 0.94 and 0.84, respectively.
@@
In HCM performances involving Khayal or Dhrupad singing, the starting 5 to 10 minutes of a performance generally involves the ”Vilambhit” (very slow tempo) section, where the singer will be singing just straight notes stretched to a few seconds, until they start singing in ”Madhya Laya” (medium tempo of the range 100 to 150). Considering that this dataset is composed of actual HCM performances, it is expected that some of the 30 second chunks from the starting portions of such singing performances will have just a few straight notes stretched for longer periods of times. These kind of audio files (30-second) will most probably confuse the model since these would come out to be very similar for any ragas which have similar ”swaras” in their ”Aaroh” and ”Avroh”. Considering that, the achieved performance is remarkable. For our current analysis, we are focused on identifying salient regions based on predictions, so we report ARI results at the chunk-level only and do not predict Raga on song-level for our dataset.
Now, we carry out further analysis in terms of ExAI analysis to understand the credibility of this classifier.

### v2

@@ IV-A Feature Extraction
For each 30-second audio chunk, we extract pitch-based chromagram features using the librosa library in Python. These features are computed with $n_{chroma}$ chromagram bins. The chromagram provides a time-frequency representation emphasizing the pitch classes, which is critical for raga analysis. However, before inputting the chromagram to the Deep Learning model, the signal undergoes tonic normalization as explained.
@@ IV-B Tonic Normalisation
Our dataset contains performances by various artists, with each performance having a single tonic note that serves as the reference note for constructing the Raga. However, different performances in the same Raga have been sung in varying tonics, leading to structural differences even within the same Raga. To address this, we perform tonic normalization on the extracted chromagram features, aligning all performances to a common reference note.
For a given audio signal $x(t)$, we first compute its Short-Time Fourier Transform (STFT), denoted as $X(n,k)$. The magnitude squared of the STFT coefficients $|X(n,k)|^{2}$ yields the spectrogram of the signal, capturing the energy distribution over time and frequency. To obtain a pitch-based log-frequency spectrogram, we map the STFT frequency bins to the 128 MIDI pitch values. The center frequency of each MIDI pitch $p$ is given by [36]:
$F_{\text{pitch}}(p)=440\times 2^{\frac{(p-69)}{12}},\quad\text{for }p\in[0,127]$ (1)
For each pitch $p$, we define a set $P(p)$ as:
$P(p)=\left\{k\mid F_{\text{pitch}}(p-0.5)\leq F_{\text{coef}}(k)<F_{\text{pitch}}(p+0.5)\right\}$ (2)
This effectively groups the STFT frequency bins into 128 pitch bands corresponding to the MIDI pitches. Now, we have the log-frequency spectrogram $Y_{\text{LF}}(n,p)$ given by:
$Y_{\text{LF}}(n,p)=\sum_{k\in P(p)}|X(n,k)|^{2}$ (3)
This reduces the frequency axis of the original spectrogram (which has $(N_{\text{FFT}}/2)+1$ frequency bins) to 128 values, each corresponding to a MIDI pitch. To obtain chromagram features, we simply sum up the pitch values which differ by some integer multiple of an octave as they are perceived to be the same note musically[36]:
@@
Now, to shift the chromagram representation of any song by a factor of ``$a"$ semitones, we are essentially looking to get a new $X^{\prime}(n,k)$ such that:
@@
Shifting by ``$a"$ semitones essentially means multiplying the frequency values in a spectrogram by $2^{a/12}$, where ``$a"$ is a discrete value such that $a\in[1:11]$. We choose $a$ within this range because shifting by 12 semitones (an octave) would map notes back to their original pitch classes, which is not desired for this transformation. Now, we will get the log-frequency spectrogram as:
@@
Equation 8 shows that shifting the frequency bins by $a$ semitones corresponds to shifting the MIDI pitch numbers by $a$ steps downward. Thus, each pitch $p$ in the shifted spectrogram $Y_{\text{LF}}^{\prime}\left(n,p\right)$ maps to pitch $p-a$ in the original spectrogram $Y_{\text{LF}}\left(n,p\right)$. This implies that shifting the chromagram by $a$ semitones is equivalent to a cyclic shift of the pitch axis by $a$ steps. Since the chromagram representation has a periodicity of 12 semitones (an octave), shifting by $a$ semitones wraps the pitch classes cyclically. So, to apply tonic normalization to our audio data, we can simply perform a cyclic shift of the extracted chromagram bins along the pitch axis by $a$ steps.
@@ IV-C Classification Model Description
To determine the best-performing model, we conduct an ablation study, experimenting with various CNN and LSTM-based architectures. Our first model, CN1+T, utilizes a simple CNN architecture with convolutional, max-pooling, and batch normalization layers, followed by a flatten layer, in a very similar way to a common image classification model. Then the CN1+LSTM+T model extends this by replacing the flatten layer with an LSTM layer, ideal for processing time-series data like audio signals. Also, we provide tonic normalization as discussed in Section:IV-B and test the model performance. The CN2+LSTM model features a more complex CNN-LSTM architecture with 1D max-pooling to preserve temporal information, as shown in Fig. 2. Then the CN2+LSTM+T model just includes tonic normalization in the input features, following the exact configuration of CN2+LSTM model. The selection of layer configurations and activation functions is carefully optimized for classification performance. These models will be further detailed in sectionV, where we evaluate their effectiveness in ARI.
@@ IV-D Classification Model Training
We train our DL models for the 12-class classification task using softmax activation in the last layer. Let $z_{i}$ be the output logit for the $i^{\text{th}}$ neuron in the last dense layer. The softmax function converts these logits into probability values:
$p_{i}=\text{softmax}(z_{i})=\frac{e^{z_{i}}}{\sum_{j=1}^{N}e^{z_{j}}}$ (9)
where $N=12$ is the number of classes. The probabilities $p_{i}$ satisfy $\sum_{i=1}^{N}p_{i}=1$. We use the categorical cross-entropy loss function for training:
$L_{\text{CCE}}=-\sum_{i=1}^{N}t_{i}\log(p_{i})$ (10)
where $t_{i}$ is the true label encoded as a one-hot vector (i.e., $t_{i}=1$ for the correct class and 00 elsewhere). Training is performed over multiple epochs until convergence, and model performance is evaluated using precision, recall, and F1-score. We employ cross-validation to assess the robustness and generalization ability of our model.
@@ IV-E Grad-CAM++
We utilize the CNN component of our trained model without the LSTM layers to get 256 learned feature maps in our case, as is visible in Fig. 2. To create a class-specific GradCAM++ activation map, each of these feature maps is weighted, and then a weighted sum is taken to generate the final activation map or saliency map for the target class. We compute the weights of feature maps using the weights of those last CNN blocks, given by the equation:
@@
Here, $A_{ij}^{k}$ represents the activation at position $(i,j)$ in the $k^{\text{th}}$ feature map. We have $k\in[1,256]$ feature maps, and $c\in[1,12]$ classes. $\alpha_{c}^{k}$ represents the weights of $k^{th}$ feature map concerning the $c^{th}$ class. $Z$ is the dimension of each feature map, and $y^{c}$ represents the output logits for class c. By computing these weights $\alpha_{c}^{k}$, we perform a weighted sum of the feature maps to obtain the GradCAM [8] attribution map given by:
@@
After that, ReLU is applied to keep only the positively impacting salient regions and remove the regions with negative gradients in the output saliency map or attribution map. Now, we weigh each pixel separately rather than weighting the whole feature map with a single weight. The new weights are given by:
@@
Here, $A^{k}_{ij}$ represents each pixel in the $k^{th}$ feature map. This $w^{kc}_{ij}$ aims to weigh each pixel separately rather than the whole activation map. So, this $w^{kc}_{ij}$ is given by:
@@
Using these, we finally obtain the attribution map for Grad-CAM++ [9] given by:
@@ IV-F SoundLIME
SoundLIME[10] is specifically designed for audio, analogous to LIME’s[33] application in images. Here also, we need the pre-trained classifier model, but we treat it as a black-box model with no access to its gradients. We temporally divide the chromagram into 30 equal parts or super-pixels, with one super-pixel representing the chromagram representation for each second of the audio. Next, we create 150 random perturbations for the given test input by applying random multi-hot masks of shape 30x1, meaning that each element of the mask either suppresses or retains each super-pixel. An exponential kernel is defined on L2 distance function $D(x,z)$ between the original image $x$ and the perturbed image $z$, denoted by $\pi_{x}(z)$:
@@ IV-G Saliency Evaluation Method
To evaluate the salient regions identified by the two XAI models, we use the output of XAI models and compare them with the expert-annotated salient regions. We compute the precision for top 1 to 10 seconds for both the XAI models. In our task, the concern lies solely with determining whether the top predicted regions are correct. Hence, precision is the best metric to employ in this situation, whereas recall and accuracy are not required. We then analyse one correct and one incorrect prediction of the classifier for their salient region.
@@ V-A Data Pre-processing
To prepare the audio files for analysis, all non-musical speech is removed. We distribute all the songs into train-set $S_{tr}$, test-set $S_{t}$, and validation-set $S_{val}$ with percentages of approximately 79%, 15%, and 6% respectively, making sure that there is at least 1 song of each class in the test and validation sets. Then we split the songs into 30-second chunks for all the 3 splits as shown in Fig. 3. Notice that we first split the dataset into training, validation, and test sets at the song level, and then divide audios in each set into 30-second chunks. This ensures that the chunks from the same song are not distributed across different sets, which would introduce bias and compromise the generalizability of our model.
@@
The last chunk of each music recording is purposely removed as it is mostly found to have only 3 to 10 seconds of music followed by silence, or may simply consist of a single ending note prolonged to a few seconds, which may not even contain any salient region for getting classified to a particular Raga.
We extract chromagram features for these audio snippets, making each input sample of shape (938,12), having 938 time frames with 12 chromagram bins each. We decide Note “A” to be the reference note for the dataset and carry out tonic normalization as explained in Section IV-B, using the tonic values labeled by the music experts. Now we carry out training using the train-set $S_{tr}$ and validation-set $S_{val}$ as explained in Section IV-C,IV-D and then evaluate on test-set $S_{t}$.
@@ V-B Classification results
Chunk-wise classification results on $S_{t}$ after carrying out ablation studies, and cross-validation are shown in Table I (all values rounded off to 2 decimal places)
TABLE I: Combined Classification Report for Different Models
@@ V-B1 CN1+T
This model achieved an overall weighted f1-score of 0.78 on $S_{t}$ for $N=12$ classes. The maximum and minimum f1-scores during cross-validation were 0.80 and 0.73, respectively. Clearly, this architecture is not suitable for music data, as flattening does not appropriately handle the dependent frame-wise information required in case of such time-series data.
@@ V-B2 CN1+LSTM+T
This model showed improved classification performance with an f1-score of 0.83 after cross-validation. The maximum and minimum f1-score observed during cross-validation were 0.87 and 0.79, respectively. This model outperformed the CN1+T model, highlighting the importance of incorporating temporal modeling layers, such as LSTM, instead of flattening layers to effectively capture sequential patterns in music-based applications.
@@ V-B3 CN2+LSTM
Despite multiple training attempts, extensive hyperparameter tuning, and architectures similar to our main model (CN2+LSTM+T) or its deeper variants with additional CNN and LSTM layers, this model failed to achieve satisfactory performance even on $S_{tr}$. As a result, its performance on the $S_{t}$ was expectedly poor, underscoring the importance of tonic normalization for the ARI task. We also experiment with simplifying the problem by carrying out training on a smaller subset of $S_{tr}$ having $N<12$ classes, removing the classes with very low f1-scores during training. While this led to some improvement in training, it suggests a lack of generalization to any number of Raga classes using this approach.
@@ V-B4 CN2+LSTM+T
This model consistently achieved the highest performance among all models, with an average f1-score of 0.89 after 6-fold cross-validation. The maximum and minimum f1-scores observed during cross-validation are 0.94 and 0.84, respectively. While the CN1+LSTM+T model showed better classification results for four classes, the proposed model closely follows, demonstrating strong training across all 12 classes.
@@
In HCM performances involving Khayal or Dhrupad singing, which constitutes the majority of our dataset, the starting 5 to 10 minutes of the performance generally involves the “Vilambhit” (very slow tempo) section, where the singer will be singing just straight notes stretched to a few seconds until they start singing in “Madhya Laya” (medium tempo of the range 100 to 150 BPM). Consequently, some of the 30-second audio clips extracted from these early portions contain only a few long, sustained notes. Such clips are likely to challenge the model, as they may appear very similar across different ragas that share similar “swaras” in their “Aaroh” (Bihag and Maru-Bihag) and “Avroh” (Shuddha-Kalyan and Yaman). Despite this inherent difficulty, the model demonstrates remarkable performance.
V-C Testing on Other Datasets
We test our CN2+LSTM+T model on the Saraga dataset[12] and test examples from YouTube. We carefully select audio files from the Saraga dataset corresponding to the 12 Raga classes in our study. We do not find any test files for Nayaki-Kanada in Saraga, so we use test files from our original dataset $S_{t}$ for this particular class. For the remaining Raga classes, we identify a total of 23 audio files from Saraga, with Bhairavi having the maximum representation of 5 files, while other classes have 3 or fewer, some with only a single file per class. From YouTube, we select 55 files of HCM performances having a minimum of 4 and a maximum of 5 files per Raga class. We apply all pre-processing as described in Section V-A. We utilize the tonic values given in the metadata of Saraga dataset. We test the compiam[37] tool on our dataset and find it to be 91.76% accurate in predicting tonic values. So, we use this to get tonic values for the YouTube test files. Our evaluation metrics include both chunk-wise f1-score and song-wise (polling among all the chunks of each song for final prediction) f1-score. The comparative results between our test dataset $S_{t}$, YouTube test data, and Saraga dataset are presented in Table II.
The performance metrics displayed in Table II are notably robust, especially considering that the model is not trained on the Saraga dataset or any of the YouTube files, but merely tested on them. These results not only demonstrate the efficacy of our proposed model but also validate the quality and representativeness of our newly curated dataset. The model’s ability to generalize well to an external dataset underscores its potential for practical applications in Raga classification tasks.
TABLE II: Test Results on Saraga and YouTube Test Examples
Dataset Chunk-level Song-level
Precision Recall F1 Precision Recall F1
$S_{t}$ 0.90 0.89 0.89 0.97 0.97 0.97
Saraga 0.82 0.76 0.78 0.92 0.88 0.90
YouTube 0.82 0.78 0.80 0.93 0.92 0.92
TABLE III: Classification report after training on the whole dataset with 41 distinct Raga classes and 42nd class being the ‘others’ class, which includes all the remaining Raga classes.
True Precision Recall F1-Score
Chunk-Level 0.65 0.64 0.64
Song-Level 0.80 0.82 0.81
V-D Train on the whole dataset
We also train the CN2+LSTM+T model on the whole dataset using the top 42 Raga classes having at least 4 audio files, and the 43rd class being the Others class, consisting of audio files from all the remaining Raga classes. The model demonstrates effective learning on the whole dataset, achieving a chunk-level f1-score of 0.64 and song-level f1-score of 0.80 as shown in Table III. While these results are not directly relevant to our XAI analysis, we report them to establish a performance benchmark for our dataset when applied across the full range of Raga classes.